In reading the instruction manual that came with your garagedoor opener, you see that the transmitter unit in your car produces a 250 mW, 350 MHz signal and that the receiver unit is supposed to respond to a radio wave at this frequency if the electric field amplitude exceeds 0.10 V/m. You wonder if this is really true. To find out, you put fresh batteries in the transmitter and start walking away from your garage while opening and closing the door. Your garage door finally fails to respond when you’re 42 m away.  What is the electric field amplitude at the receiver when it first fails?
Answer #1

formula for intensity

I = P/A = P/ 4pi r^2 = 0.250/ 4pi ( 42)^2 = 1.128 x 10^-5 W/m^2

I = E o^2/ 2 uo c

E o = sqrt 2 uo c I

= sqrt 2 (4 pi * 10^-7 (3 * 10^8) ( 1.128 x 10^-5 W/m^2)

=0.0922 V/m